# chebgreen—: Learning and Interpolating Continuous Empirical Green’s Functions from Data

### v2

@@ 3 Numerical results
  • Generating the dataset: The length scale of the squared exponential covariance kernel in the Gaussian Process from which the forcing terms are sampled is set to $\sigma=10^{-2}$. The forcing terms and system responses are measured on a uniform 1D grid at $N_{\text{x}}=N_{\text{s}}=500$ equally spaced points. It is crucial that the density of sampling points is chosen so that functions are sufficiently resolved on the grid, i.e., and the discrete vector of values at the sampling locations are free from spatial aliasing. In our case, since the forcing terms and responses are sampled from a continuous representation in a Chebyshev basis (using chebfun), we use the number of Chebyshev points used to represent these functions as a proxy for the number of samples needed on a uniform grid. Note that equispaced points suffer from Runge Phenomenon [45] so a uniform grid is not optimal for representing these functions but data are generally collected at uniform grids in experiments; thus we chose to test our methods in a such a setting. Finally, the number of input-output pairs is set to $N_{\text{samples}}=250$, out of which only ${}95\%$ are used during training.
@@
  • chebpy2: When learning polynomial approximations for one and two-dimensional functions using chebpy2, we specify the accuracy of the functions that need to be resolved in the domain as $\varepsilon^{\text{cheb}}\approx 2.22\times 10^{-16}$. In case of two dimensions, we specify this value for each of the independent variable within the domain, $\varepsilon^{\text{cheb}}_{x}$ and $\varepsilon^{\text{cheb}}_{s}$ for $G(x,s)$. These are both set to $\varepsilon^{\text{cheb}}_{x}=\varepsilon^{\text{cheb}}_{s}\approx 2.22\times 10^{-16}$. We choose the value of $\approx 2.22\times 10^{-16}$, which is the eps - the difference between $1.0$ and the next smallest representable floating point number larger than $1.0$ for single floating point precision, in a direction for all experiments as our neural networks are trained with float64 (double floating point) precision. In the cases where there is noise in the collected data, as is the case when data are collected from experiments, this number can be set close to the noise floor so that the model does not overfit to noise. However, in our numerical experiments with noise, we do not change this number, so as to demonstrate the robustness of our method to noise.
@@
[…]
@@ 3.1 Poisson problem
Before we demonstrate our method’s efficacy for interpolation, we will approximate the Green’s function, and it’s singular value expansion, for a one-dimensional Laplacian operator. Assume we have a Poisson problem with homogeneous boundary conditions defined as follows:
$-\frac{d^{2}}{dx^{2}}u(x)=f(x),\quad x\in[0,1],\quad u(0)=u(1)=0.$ (18)
For this is a canonical problem, the closed-form solution for the associated Green’s function is known and is given by:
@@
as an error contour. The relative error $\epsilon=0.36\%$ indicates satisfactory the fidelity of our learned Green’s function to the analytical solution. Finally, we observe that the first five learned left (Fig. 8D) and right (Fig. 8F) singular functions are indistinguishable from the left and right singular functions of the analytical solution of the Green’s function. In Fig. 8E, we plot the first 100 learned singular values for the Green’s function against the analytical values of the singular values, $1/(\pi^{2}k^{2})$. There is an exponential decay for the smallest singular values in case of the learned Green’s function. We use Rational NN to approximate the Green’s function which is a smooth approximation to the exact Green’s function; thus, this fast decay is expected. The key point is that the largest learned singular values are in close agreement with analytical solution. Therefore, our method constructs a high-fidelity low-rank approximation of the Green’s function associated with Laplacian operator.
@@
Fig. 9A shows how the relative error for the learned Green’s function, $\epsilon$, changes as a function of the noise level, $\zeta$, for the system responses in the training dataset. Note that we only consider the effect of output noise in this case. An instance of a possible sample from the dataset with $\zeta=0.6$ ($60\%$ noise) is shown in Fig. 9: B is the forcing function used to perturb the system, and C is the corresponding clean system response (blue) along with the artificially polluted system response (red). Finally, the error contour of the Green’s function learned with a dataset with noise level $\zeta=0.6$ against the analytical Green’s function is shown in Fig. 9D. The relative error is $\epsilon=1.38\%$. Thus, even when the system responses barely resemble the underlying function (because of noise), our method seems to works well; learning a reasonably accurate approximation to the underlying Green’s function.
@@ 3.2 Advection-Diffusion
Figure 10: (A-C) Approximation of the Green’s functions associated with the Advection-Diffusion problem at $\theta=1,2,3$ used by the interpolation scheme. (D) Interpolated Green’s function at $\theta_{*}=2.5$. (E) The target Green’s function approximated by chebgreen—.
@@ 3.3.1 Interpolation
For the interpolation case, we compute the three approximations to the Green’s functions for the parameterized operator at $\theta_{1}=1.0$, $\theta_{2}=5.0$, and $\theta_{3}=10.0$ (we call these the ”interpolant” cases). The left and right singular functions for this problem vary significantly as we change the value of $\theta$ as seen in Fig. 11, which makes it a useful test for our interpolation method.
@@
Note that since the values of the parameter $\theta$ are much further apart compared to the Advection-Diffusion case, and $\theta$ appear as a squared quantities in the operator, we expect the differences in the learned Green’s function to be much more drastic. This is evident from the learned Green’s function shown in Fig. 12A-C. The interpolated Green’s function at $\theta_{*}$ has a test error of $\epsilon_{\text{test}}=2.76\%$. This is larger than the test errors for the interpolant cases (less than $0.69\%$). Although given we are interpolating using only three points on an non-linear manifold, our method seems to perform reasonably well.
@@
Figure 12: (A-C) Approximation of the Green’s functions associated with the Airy problem at $\theta=1,5,10$ used by the interpolation scheme. (D) Interpolated Green’s function at $\theta_{*}=7$. (E) The target Green’s function approximated by chebgreen—.
@@ 3.3.2 Extrapolation
In order to demonstrate the ability of our method to generalize out of the neighborhood of the approximated Green’s functions, we compute chebgreen approximations at $\theta_{1}=6.0$, $\theta_{2}=7.0$, and $\theta_{3}=8.0$ and extrapolate to $\theta_{*}=9.0$. This satisfactory performance is due to our having learned a suitable solution operator (i.e., Green’s function). A visual comparison between the extrapolated model and a target model constructed with data at $\theta_{*}=9.0$ is shown in Fig. 13.
@@
Figure 13: (A) Extrapolated Green’s function. (B) The target Green’s function approximated by chebgreen— at $\theta_{*}=9$.
@@ 3.4 Fractional Laplacian
where $10<\theta<$ 1 is the fractional order. To generate the dataset for this problem, we use a Fourier spectral collocation to solve Eq. 19. The forcing and system responses are sampled at a slightly smaller uniform 1D grid, with $N_{x}=N_{s}=600$ samples. For chebfun, we set $\varepsilon_{x}^{\text{cheb}}=\varepsilon_{s}^{\text{cheb}}=1\times 10^{-9}$ as the Green’s function tends to be non-differentiable throughout the domain which makes it difficult to resolve this to float64 precision.
@@
Figure 14: (A-C) Approximation of the Green’s functions associated with the Fractional Laplacian problem at $\theta=0.8,0.9,0.95$ used by the interpolation scheme. (D) Interpolated Green’s function at $\theta_{*}=0.85$. (E) The target Green’s function approximated by chebgreen—.
3.5 Summary of errors for all problems
The test errors for all the interpolant, interpolated, and target Green’s functions are summarized in Table 1. We also compute the test errors when the dataset has $50\%$ output noise (noise level $\zeta=0.5$). Note that here we learn the Green’s function with the noisy dataset but we test with a clean dataset; to demonstrate that we indeed learn the underlying solution operator. The test error on the noisy dataset for these cases are available in Table 2.
Problem $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Interpolated Target
Airy 0.38 (1.87) 0.69 (1.91) 0.62 (2.02) 2.76 (3.28) 0.76 (1.84)
Fractional Laplacian 0.51 (2.38) 0.95 (2.81) 0.65 (1.87) 0.99 (2.94) 0.41 (2.21)
Advection Diffusion 0.24 (1.57) 0.26 (1.49) 0.29 (1.49) 0.54 (1.61) 0.3 (1.54)
Table 1: Summary of the test error for the Airy, Fractional Laplacian, and Advection-Diffusion problems. The error for experiments with noisy dataset are shown in parenthesis.
@@ 4 Conclusions
In this work, we propose an extension of the previous work [30] on learning and interpolating Green’s function from data. Some notable improvements include the inclusion of greenlearning [29] to learn an approximation of Green’s function in form of a continuous, bivariate function and the use of our python implementation of chebfun library [33], chebpy, to learn a Singular Value Expansion instead of Singular Value Decomposition, in the previous work [30]. This allows the new method to learn a representation which is not tied to a particular discretization of the domain, making it truly mesh-independent. Finally, the novel generalization of the previous interpolation algorithm for orthonormal matrices on Stiefel manifold to the interpolation of Quasimatrices on the infinite dimensional analogue of Stiefel manifold, provides a way to seamlessly use the machinery we have created in order learn the solution operator for 1D parametric linear partial differential equations. We present multiple examples in order to demonstrate that the method learns high-fidelity approximations to the underlying Green’s function of a linear operator with an order of magnitude lower data requirement compared to the previous work. We also demonstrate that the proposed method is robust against large amounts of noise in the dataset. This, along with the easy to use library which is available online makes it a useful tool for use in discovering the solution operators in realistic experimental settings.
Data availability
The code used to produce the numerical results is publicly available on GitHub at https://github.com/hsharsh/chebgreen for reproducibility purposes.
Acknowledgments
This work was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Numbers N00014-23-1-2729 and N00014-22-1-2055.
Appendix A Proof for interpolating quasimatrices on tangent space of $H=(L^{2}(\Omega))^{K}$ space
Theorem 1.
The map $P_{\Phi}\colon H\rightarrow T_{\Phi}S_{K}(H)$ given by $\Psi\mapsto\Psi-\Phi\operatorname{sym}(\Phi^{*}\Psi)$ is the orthogonal projection onto the tangent space of the Stiefel manifold $S_{K}(H)$ at $\Phi,$ where $\operatorname{sym}(\Phi^{*}\Psi)\coloneq\frac{1}{2}(\Phi^{*}\Psi+\Psi^{*}\Phi).$
@@ Appendix A Proof for interpolating quasimatrices on tangent space of $H=(L^{2}(\Omega))^{K}$ space
Let $A=\Phi^{*}\Psi$ and $B=\Psi^{*}\Phi.$ First we show that $P_{\Phi}$ is a projection operator, i.e. that $P_{\Phi}\circ P_{\Phi}=P_{\Phi}.$ For $\Psi\in H,$ we have
$P_{\Phi}(P_{\Phi}(\Psi))=\Psi-\frac{1}{2}\Phi(A+B)-\frac{1}{2}\Phi\left([\Psi-\frac{1}{2}\Phi(A+B)]^{*}\Phi+\Phi^{*}[\Psi-\frac{1}{2}\Phi(A+B)]\right)$
$=\Psi-\frac{1}{2}\Phi(A+B)-\frac{1}{2}\Phi\left(\Psi^{*}\Phi-[\frac{1}{2}\Phi(A+B)]^{*}\Phi+\Phi^{*}\Psi-\Phi^{*}[\frac{1}{2}\Phi(A+B)]\right)$
$=\Psi-\frac{1}{2}\Phi(A+B)-\frac{1}{2}\Phi\left(A-\frac{(A+B)}{2}\Phi^{*}\Phi+B-\Phi^{*}\Phi\frac{(A+B)}{2}\right)$
Since $\Phi$ is an element of the Stiefel manifold, it satisfies $\Phi^{*}\Phi=I_{K}.$ So all of the terms in the last line to the right of $\Psi-\frac{1}{2}\Phi(A+B)$ cancel out, and it follows that $P_{\Phi}P_{\Phi}=P_{\Phi}$ as desired.
Next, we must show that $P_{\Phi}$ is self-adjoint in order to establish that it is an orthogonal projection. For any $\Psi,\Gamma\in H,$ we have
$\operatorname{tr}\left(P_{\Phi}(\Psi)^{*}\Gamma\right)=\operatorname{tr}\left(\Psi^{*}\Gamma-\frac{1}{2}\left[\Phi(\Phi^{*}\Psi+\Psi^{*}\Phi)\right]^{*}\Gamma\right),$

### v5

@@ 4 Numerical results
  • Generating the dataset: The length scale of the squared exponential covariance kernel in the Gaussian Process from which the forcing terms are sampled is set to $\sigma=10^{-2}$. The forcing terms and system responses are measured on a uniform 1D grid at $N_{\text{x}}=N_{\text{s}}=500$ equally spaced points. It is crucial that the density of sampling points is chosen so that functions are sufficiently resolved on the grid, i.e., and the discrete vector of values at the sampling locations are free from spatial aliasing. In our case, since the forcing terms and responses are sampled from a continuous representation in a Chebyshev basis (using chebfun), we use the number of Chebyshev points used to represent these functions as a proxy for the number of samples needed on a uniform grid. Note that equispaced points suffer from Runge Phenomenon [Trefethen_Runge] so a uniform grid is not optimal for representing these functions but data are generally collected at uniform grids in experiments; thus we chose to test our methods in a such a setting. Finally, the number of input-output pairs is set to $N_{\text{samples}}=100$, out of which only $95\%$ are used during training. The choice of $100$ samples is motivated by a study on the fidelity in learning the Green’s function for the Laplacian operator in Section 4.1.1.
@@
  • chebpy2: When learning polynomial approximations for one and two-dimensional functions using chebpy2, we specify the accuracy of the functions that need to be resolved in the domain as $\varepsilon^{\text{cheb}}\approx 2.22\times 10^{-16}$. In case of two dimensions, we specify this value for each of the independent variable within the domain, $\varepsilon^{\text{cheb}}_{x}$ and $\varepsilon^{\text{cheb}}_{s}$ for $G(x,s)$. These are both set to $\varepsilon^{\text{cheb}}_{x}=\varepsilon^{\text{cheb}}_{s}\approx 2.22\times 10^{-16}$. We choose the value of $\approx 2.22\times 10^{-16}$, which is the eps - the difference between $1.0$ and the next smallest representable floating point number larger than $1.0$ for single floating point precision, in a direction for all experiments as our neural networks are trained with float64 (double floating point) precision. This choice affects the rank $K$ of the learned Green’s function approximation. The rank $K$, defined as the number of rank $1$ functions in the bivariate representation in Eq. 10, is determined adaptively and depends on the specified error tolerances; $K$ increases as $\varepsilon^{\text{cheb}}_{x}$ and $\varepsilon^{\text{cheb}}_{s}$ are reduced for finer resolution. In the cases where there is noise in the collected data, as is the case when data are collected from experiments, this number can be set close to the noise floor so that the model does not overfit to noise. However, in our numerical experiments with noise, we do not change this number, so as to demonstrate the robustness of our method to noise.
@@
$\epsilon_{\text{test}}=\dfrac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\dfrac{\|\tilde{u}_{i}-u_{i}\|_{L^{2}(\Omega)}}{\|u_{i}\|_{L^{2}(\Omega)}}\times 100\%,\quad\tilde{u}_{i}=\int_{\Omega}\mathcal{G}(x,s)f_{i}(s)ds+\mathcal{H}(x).$ (17)
@@ 4.1 Poisson problem
Before we demonstrate our method’s efficacy for interpolation, we will approximate the Green’s function, and its singular value expansion, for a one-dimensional Laplacian operator. Assume we have a Poisson problem with homogeneous boundary conditions defined as follows:
$\frac{d^{2}}{dx^{2}}u(x)=f(x),\quad x\in[0,1],\quad u(0)=u(1)=0.$ (18)
For this is canonical problem, the closed-form solution for the associated Green’s function is known and is given by:
@@
as an error contour. The relative error $\epsilon=0.36\%$ indicates satisfactory the fidelity of our learned Green’s function to the analytical solution. Finally, we observe that the first five learned left (Fig. 8D) and right (Fig. 8F) singular functions are indistinguishable from the left and right singular functions of the analytical solution of the Green’s function. In Fig. 8E, we plot the first 100 learned singular values for the Green’s function against the analytical values of the singular values, $1/(\pi^{2}k^{2})$. There is an exponential decay for the smallest singular values in the case of the learned Green’s function. We use Rational NN to approximate the Green’s function which is a smooth approximation to the exact Green’s function; thus, this fast decay is expected. The key point is that the largest learned singular values are in close agreement with the analytical solution. Therefore, our method constructs a high-fidelity low-rank approximation of the Green’s function associated with the Laplacian operator.
@@
4.1.1 Study on effects of number of samples, $N_{\text{samples}}$
To determine an appropriate number of samples, $N_{\text{samples}}$, for achieving sufficient fidelity to the solution operator, we conducted a study examining the relationship between the number of samples used and the resulting error. Our analysis revealed a significant reduction in error as the sample count increased, initially. However, this improvement exhibited diminishing returns, plateauing at approximately $100$ samples. As the theoretical results [Boulle2021_Theory] suggest, for a covariance kernel of a given quality, only a relatively small degree of oversampling is necessary to learn a rank $K$ approximation to underlying Green’s function. Consequently, the plateau in error reduction observed around $100$ samples likely indicates that we have extracted the majority of the learnable information about the underlying solution operator, given the characteristics of our GP kernel and the quality of our training data. Therefore, we proceeded with 100 samples in our subsequent numerical experiments, balancing computational cost with achieving a level of fidelity consistent with the theoretical limits.
[…]
Figure 9: (A) Relative error for the Laplacian operator when trained with different number of samples, $N_{\text{train}}$, (B) Contours for relative error between the learned model with a noisy dataset at $N_{\text{train}}=128$ and analytical solution for the Laplacian operator.
4.1.2 Study on effects of noise in the samples
Fig. 10A shows how the relative error for the learned Green’s function, $\epsilon$, changes as a function of the noise level, $\zeta$, for the system responses in the training dataset. Note that we only consider the effect of output noise in this case. An instance of a possible sample from the dataset with $\zeta=0.6$ ($60\%$ noise) is shown in Fig. 10: B is the forcing function used to perturb the system, and C is the corresponding clean system response (blue) along with the artificially polluted system response (red). Finally, the error contour of the Green’s function learned with a dataset with noise level $\zeta=0.6$ against the analytical Green’s function is shown in Fig. 10D. The relative error is $\epsilon=1.29\%$. Thus, even when the system responses barely resemble the underlying function (because of noise), our method seems to works well; learning a reasonably accurate approximation to the underlying Green’s function.
@@ 4.2 Advection-Diffusion
Figure 11: (A-C) Approximation of the Green’s functions associated with the Advection-Diffusion problem at $\theta=1,2,3$ used by the interpolation scheme. (D) Interpolated Green’s function at $\theta_{*}=2.5$ ($\epsilon_{\text{test}}=0.54\%$). (E) The target Green’s function approximated by chebgreen—.
@@
[…]
Figure 12: Relative error at different noise level, $\zeta$, for the interpolated Advection-Diffusion operator at $\theta=2.5$.
We also show that our interpolated model is robust to noise in the training data used to create the individual interpolants (Fig. 12). This figure illustrates how the test error ($\epsilon_{\text{test}}$) of the interpolated Green’s function changes as the output noise level ($\zeta$) increases in the system response data used to train the interpolants. Note that even when the level of noise in the training data is $\zeta=1.0$ ($100\%$), the test error is still less than $5\%$.
@@ 4.3.1 Interpolation
For the interpolation case, we compute the three approximations to the Green’s functions for the parameterized operator at $\theta_{1}=1.0$, $\theta_{2}=5.0$, and $\theta_{3}=10.0$ (we call these the “interpolant” cases). The left and right singular functions for this problem vary significantly as we change the value of $\theta$ as seen in Fig. 13, which makes it a useful test for our interpolation method.
@@
Note that since the values of the parameter $\theta$ are much further apart compared to the Advection-Diffusion case, and $\theta$ appear as a squared quantities in the operator, we expect the differences in the learned Green’s function to be much more drastic. This is evident from the learned Green’s function shown in Fig. 14A-C. The interpolated Green’s function at $\theta_{*}$ has a test error of $\epsilon_{\text{test}}=2.76\%$. This is larger than the test errors for the previously discussed interpolant cases (less than $0.69\%$). Although, given we are interpolating using only three points on a non-linear manifold, our method seems to perform reasonably well.
@@
Figure 14: (A-C) Approximation of the Green’s functions associated with the Airy problem at $\theta=1,5,10$ used by the interpolation scheme. (D) Interpolated Green’s function at $\theta_{*}=7$ ($\epsilon_{\text{test}}=2.76\%$). (E) The target Green’s function approximated by chebgreen—.
@@ 4.3.2 Extrapolation
In order to demonstrate the ability of our method to generalize out of the neighborhood of the approximated Green’s functions, we compute chebgreen approximations at $\theta_{1}=6.0$, $\theta_{2}=7.0$, and $\theta_{3}=8.0$ and extrapolate to $\theta_{*}=9.0$. This satisfactory performance is due to our having learned the governing solution operator (i.e., Green’s function). A visual comparison between the extrapolated model and a target model constructed with data at $\theta_{*}=9.0$ is shown in Fig. 15. The extrapolated Green’s function at $\theta_{*}$ has a test error of $\epsilon_{\text{test}}=2.28\%$.
@@
Figure 15: (A) Extrapolated Green’s function ($\epsilon_{\text{test}}=2.28\%$). (B) The target Green’s function approximated by chebgreen— at $\theta_{*}=9$.
@@ 4.4 Fractional Laplacian
where $0<\theta<1$ is the fractional order. To generate the dataset for this problem, we use a Fourier spectral collocation to solve Eq. 19. The forcing and system responses are sampled at a slightly smaller uniform 1D grid, with $N_{x}=N_{s}=600$ samples. For chebfun, we set $\varepsilon_{x}^{\text{cheb}}=\varepsilon_{s}^{\text{cheb}}=1\times 10^{-9}$ as the Green’s function tends to be non-differentiable throughout the domain which makes it difficult to resolve this to float64 precision.
@@
Figure 16: (A-C) Approximation of the Green’s functions associated with the Fractional Laplacian problem at $\theta=0.8,0.9,0.95$ used by the interpolation scheme. (D) Interpolated Green’s function at $\theta_{*}=0.85$ $\epsilon_{\text{test}}=0.99\%$. (E) The target Green’s function approximated by chebgreen—.
@@ 5 Conclusions
In this work, we propose an extension of the previous work [Praveen2023_EGF] on learning and interpolating Green’s function from data. Some notable improvements include the inclusion of greenlearning [Boulle2021_Greenlearning] to learn an approximation of Green’s function in form of a continuous, bivariate function and the use of our python implementation of chebfun library [Chebfun], chebpy, to learn a Singular Value Expansion instead of Singular Value Decomposition, in the previous work [Praveen2023_EGF]. This allows the new method to learn a representation which is not tied to a particular discretization of the domain, making it truly mesh-independent. Finally, the novel generalization of the previous interpolation algorithm for orthonormal matrices on Stiefel manifold to the interpolation of Quasimatrices on the infinite dimensional analogue of Stiefel manifold, provides a way to seamlessly use the machinery we have created in order learn the solution operator for 1D parametric linear partial differential equations. We present multiple examples in order to demonstrate that the method learns high-fidelity approximations to the underlying Green’s function of a linear operator with an order of magnitude lower data requirement compared to the previous work. We also demonstrate that the proposed method is robust against large amounts of noise in the dataset. This, along with the easy to use library which is available online makes it a useful tool for use in discovering the solution operators in realistic experimental settings.
[…]
A primary direction for future investigation is the extension of our method to higher-dimensional problems. This requires careful consideration, as Eckart-Young-Mirsky theorem is not valid for tensors of order greater than two. One strategy would be using the same approach as the chebfun3| [Hashemi2017_chebfun3] framework: Using the functional Tucker representation [Dolgov2021_TuckerFuncApprox] to create continuous approximations of the Green’s functions with a Chebyshev basis, followed by the continuous analogue of the Tensor Train decomposition [Gorodetsky2019_TensorTrainContinuous] to construct a representation we can interpolate with our proposed manifold interpolation method. The other strategy would be to construct a reduced order model using Deep Learning [Brivio2024_PODDeep].
[…]
for reproducibility purposes.
[…]
$\|v-P_{\phi}(y-\phi)\|_{H}\leq C\,\|y-\phi\|_{H}^{3}.$
@@
Let $y=\exp_{\phi}(v)$ with $v\in U\subset T_{\phi}S_{K}$. We know that $U$ is within the radius of injectivity because the metric on $S_{K}$ is induced by $H$, and hence by the Taylor expansion of $\exp_{\phi}$ in normal coordinates, which exist in $U$ since $U$ is contained within the radius of injectivity of the exponential map, the first derivative is the identity and the second derivative vanishes [Boumal_2023]. Thus
$\exp_{\phi}(v)=\phi+v+\frac{1}{6}\,D^{3}\exp_{\phi}(0)[v,v,v]+r(v),$
with a remainder satisfying
$\|r(v)\|_{H}=O(\|v\|_{H}^{4}).$
Subtracting $\phi$ from both sides gives
$y-\phi=v+\frac{1}{6}\,D^{3}\exp_{\phi}(0)[v,v,v]+r(v).$
Since $v\in T_{\phi}S_{K}$ and $P_{\phi}$ is the orthogonal projection onto $T_{\phi}S_{K}$ (with $P_{\phi}(w)=w$ for all $w\in T_{\phi}S_{K}$), it follows that
$P_{\phi}(y-\phi)=v+P_{\phi}\Bigl{(}\frac{1}{6}\,D^{3}\exp_{\phi}(0)[v,v,v]+r(v)\Bigr{)}.$
By the non-expansivity of $P_{\phi}$, we have
$\|P_{\phi}(y-\phi)-v\|_{H}\leq\frac{1}{6}\,\|D^{3}\exp_{\phi}(0)[v,v,v]\|_{H}+\|r(v)\|_{H}.$
Since $D^{3}\exp_{\phi}(0)$ is a bounded trilinear map on $U$, there is a constant $C_{3}>0$ such that
$\|D^{3}\exp_{\phi}(0)[v,v,v]\|_{H}\leq C_{3}\,\|v\|_{H}^{3}.$
Thus,
$\|P_{\phi}(y-\phi)-v\|_{H}\leq\frac{C_{3}}{6}\,\|v\|_{H}^{3}+O(\|v\|_{H}^{4}).$
Finally, because by the local rigidity of $\exp_{\phi}$ we have
$\|y-\phi\|_{H}=\|v\|_{H}+O(\|v\|_{H}^{3}),$
it follows (after possibly adjusting the constant) that
$\|P_{\phi}(y-\phi)-v\|_{H}=O(\|v\|_{H}^{3})=O(\|y-\phi\|_{H}^{3}).$
Proposition 2.
Let $M\subset H$ be an embedded Hilbert manifold in a Hilbert space $H$, so that each tangent space $T_{\phi}M$ is a linear subspace of $H$. Suppose that
$R:TM\to M,\quad(\phi,\eta)\mapsto R(\phi,\eta)$
is a $C^{2}$ retraction, that is that:
  1. (i)
    $R(\phi,0)=\phi$.
  2. (ii)
    $DR(\phi,0)[\eta]=\eta$ for all $\eta\in T_{\phi}M$.
  3. (iii)
    $R(\phi,\cdot)$ is $C^{2}$. Hence, by continuity there exists a constant $L>0$ (depending on $\epsilon$) such that
    $\|D^{2}R(x,0)[\eta,\eta]\|_{H}\leq L\,\|\eta\|_{H}^{2}\quad\text{for all }\eta\in B_{\epsilon}(0).$
Then for every $x\in M$ and every $\eta\in T_{x}M$ with $\|\eta\|_{H}<\epsilon$, there exists $\delta>0$ (with $\delta\leq\epsilon/\|\eta\|_{H}$) such that for all $t$ with $|t|<\delta$ we have
$\|R(x,t\eta)-(x+t\eta)\|_{H}\leq L\,t^{2}\,\|\eta\|_{H}^{2}.$
Proof.
Since $R(\phi,\cdot):T_{\phi}M\to M$ is smooth on $B_{\epsilon}(0)$, Taylor’s theorem in the Banach space $T_{\{}\phi\}M$ (with norm $\|\cdot\|_{H}$ implies that for each $\phi\in M$ and each $\eta\in T_{\phi}M$ with $\|\eta\|_{H}<\epsilon$ there exists a remainder term $r(t\eta)$ with
$R(\phi,t\eta)=R(\phi,0)+t\,DR(\phi,0)[\eta]+\frac{t^{2}}{2}\,D^{2}R(\phi,0)[\eta,\eta]+r(t\eta),$